OpenStudy (anonymous):

A cellular phone company charges a $27.95 monthly fee and $.12/min for local calls. Another company charges $12.95 a month and $.32/min for local calls. For what number of minutes of local calls are the costs of the plans the same?

OpenStudy (harsha19111999):

27.95+x(0.12)=12.95+x(0.32)

OpenStudy (harsha19111999):

=>x=75 minutes

OpenStudy (anonymous):

27.95+.12x=12.95+.32x .12x-.32x=12.95-27.95 -.2x=-15 x=-15/-.2 I hope it helps